# ]https://zf-wei.github.io

### v1

@@ I Introduction
Our findings indicate that incremental selection of nodes and subsequent removal of their adjacent edges consistently reduce ECS. However, this impact tends to be more noticeable across graph embedding methods, but node2vec and LLE, and varies across network sizes, the initial community structure strength, and the perturbation type. This aligns with the expectation that removing a greater number of edge results in a more substantial disruption of the community structure. Notably, targeted node selection leads to a more pronounced and accelerated decline in ECS compared to random node selection. This is rationalized by the fact that adversarial attacks tend to swiftly dismantle the network’s community structure. Among the seven network embedding methods considered, node2vec and LLE consistently demonstrate to be more robust for community detection within their respective graph embedding families. Overall, node2vec is the most robust graph embedding method for community detection. Our results generally concur with those of Kojaku et al. [16], which found that node2vec learns more consistent community structures with networks of heterogeneous degree distributions and size.
@@ II.2 Edge removal procedure
Second, we discuss targeted edge removal. Targeted edge removal is based on the betweenness centrality of nodes. Betweenness centrality measures the importance of a node in a network based on how often it lies on the shortest paths between other nodes [46]. Node with high betweenness centrality usually indicate nodes of vital importance to the network, so removing edges adjacent to these nodes might cripple the functionality of the network [47]. This approach models an adversarial attack in real-world situations [48, 49]. To implement edge removal based on targeted node selection, we first rank the nodes in the network by their betweenness centrality. The node with the lowest betweenness centrality has rank $1$, and the node with the highest betweenness centrality has rank $N$, where $N$ is the number of nodes in the network. We then independently generate $50$ sets of nodes, each containing $p\%$ of nodes in the network. The nodes in each set are selected by randomly sampling nodes with probability proportional to their betweenness centrality rank. We ensure that the removal of edges adjacent to each set of nodes does not disconnect the remaining part of the network, which is required by graph embedding methods. In the event that the remaining part of the network becomes disconnected after the removal of a set of edges, we omit the set of nodes and proceed with re-sampling. After each removal, we measure the similarity between the community structure of remaining network in the embedding space computed through data clustering (i.e., $k$-means; see more details in Sec. II.3) and the ground-truth community structure of the original LFR network using ECS. The overall similarity score for a given $p\%$ is calculated by averaging community similarity scores across all $50$ sets.
In our experiments, the proportion of selected nodes starts from $5\%$ with an increment of $5\%$. The upper bound of proportion depends on the network structure and the node selection method. For example, when we select a network with $N=1,000$, $\mu=0.01$, and use edge removal based on targeted node selection, it is difficult to keep the network connected when deleting edges adjacent to $35\%$ of nodes. In these cases, we use a proportion of selected nodes that is below the threshold, e.g., $5\%$, $10\%$, $15\%$, $20\%$, $25\%$, and $30\%$. In addition, we also perform data clustering on the embeddings of the original LFR networks with no edges removed and compare the results with the ground-truth community structure. This result will be recorded as “0% of nodes selected.” We also capture the variation of the measurements by computing standard deviations. As the standard deviations are negligible with respect to the mean, we do not show those results here.
@@ II.4 Community similarity
To measure the similarity between community partitions, we employ ECS [35], computed with the CluSim package [61]. Compared to other clustering similarity metrics such as the normalized mutual information (NMI) [62], ECS is a most robust metric for comparing clusters and better addresses challenges such as biases in randomized membership, skewed cluster sizes, and the problem of matching [35]. We use default parameter values in ECS computation.
ECS yield values in the range from 00 to $1$, where a higher value indicates a greater similarity between partitions.
@@ III Results and Discussion
In the subsequent line figures, unless otherwise stated, subfigures in the top row, i.e., (a) and (b), correspond to random node selection. More specifically, subfigure (a) corresponds to $16$-dimensional embeddings and subfigure (b) depicts results obtained with $32$-dimensional embeddings. Similarly, subfigures in the bottom row, i.e., (c) and (d), correspond to targeted node selection. More specifically, subfigure (c) corresponds to $16$-dimensional embeddings and subfigure (d) depicts results obtained with $32$-dimensional embeddings. We focus only in these two dimensions as smaller embedding dimensions seem to be enough to provide an accurate low-dimensional vector representation of networks [15, 16]. Recall that a higher ECS value indicates a greater similarity between network community partitions obtained from the LFR benchmark graphs and the embeddings. Thus, when analyzing line figures, higher ECS values and therefore higher positions of line curves indicate a more robust graph embedding method for community detection.
@@
Finally, Sec. III.4 presents experimental results to evaluate the robustness of our community detection methods on two real-world networks with labeled communities: the email-EU-core network [63] and the AS network [7]. These networks exhibit similar characteristics to synthetic counterparts, specifically power-law degree and community size distributions as observed in LFR benchmark graphs. Our experimental methodology and insights for real-world networks mirrors previous findings for synthetic counterparts.
@@ III.1 Matrix factorization methods
Matrix-based methods project network nodes onto a Euclidean space via eigenvectors following a similar idea as in spectral clustering [64]. In this family of methods, we include Laplacian eigenmap (LE) [1, 27], locally linear embedding (LLE) [14, 28], higher-order preserving embedding (HOPE) [1, 29], and modularized non-negative matrix factorization (M-NMF) [14, 19]. We now detail results in each of them.
@@ III.1.1 Laplacian eigenmap (LE)
In Figs. 2 and 3, the impact of edge removal based on different node selection strategy is demonstrated. The effects of the initial community partition’s strength (with stronger partitions experiencing less perturbation impact, i.e., lower $\mu$ values) and the type of perturbation (targeted node selection having a significant influence on clusters) are observed. Specifically, for networks with strong community partition (i.e., $\mu=0.01$ or $\mu=0.1$), we notice that embedding dimension tends to have a stronger effect on the decay rate of similarity, despite node selection strategy. For instance, in Fig. 2(a), with an embedding dimension of $16$, the curve corresponding to $\mu=0.1$ remains consistently stable around an ECS value of approximately $1.0$ until $60\%$ of nodes are selected. Subsequently, between $60\%$ and $85\%$ of nodes selected, the curve gradually and steadily decreases from $1.0$ to $0.84$ (i.e., $16\%$ decrease). In contrast, depicted in Fig. 2(b) with an embedding dimension of $32$, the curve representing $\mu=0.1$ remains stable around $1.0$ only until $25\%$ of node are selected. Then, between $25\%$ and $85\%$ of nodes selected, the curve decreases more sharply from $1.0$ to $0.53$ (i.e., $47\%$ decrease).
Fig. 4 helps us to better understand the effect of embedding dimension and network size. We notice that for strongly clustered networks (i.e., $\mu=0.01$), a lower embedding dimension produces higher ECS. On the contrary, for weakly clustered networks (i.e., $\mu=0.5$), lower embedding dimensions produce higher ECS for networks with small sizes and higher embedding dimensions result in higher ECS for networks with larger sizes. In both scenarios represented by Fig. 4(a) and Fig. 4(b), it is observed that, for a given fixed embedding dimension, the data clustering method employing LE yields higher ECS scores for larger networks.
Figure 2: Mean ECS using LE method over the percentage of selected nodes on LFR benchmark graphs with $1,000$ nodes. (a) and (b) use random node selection while (c) and (d) use targeted node selection.
Figure 3: Mean ECS using LE method over the percentage of selected nodes on LFR benchmark graphs with $10,000$ nodes. (a) and (b) use random node selection while (c) and (d) use targeted node selection.
@@ III.1.2 Locally linear embedding (LLE)
Figure 5: Mean ECS using LLE method over the percentage of selected nodes on LFR benchmark graphs with $1,000$ nodes. (a) and (b) use random node selection while (c) and (d) use targeted node selection.
The insights drawn from Figs. 5, 6 and 7 for LLE align closely with the insights derived for LE.
Figure 6: Mean ECS using LLE method over the percentage of selected nodes on LFR benchmark graphs with $10,000$ nodes. (a) and (b) use random node selection while (c) and (d) use targeted node selection.
@@ III.1.3 Higher-order preserving embedding (HOPE)
Higher-order preserving embedding (HOPE) is designed to maintain the similarity between nodes [1, 29]. The objective function being optimized is $\bigl{|}\mathbf{S}-\mathbf{x}\mathbf{x}^{T}\bigr{|}$, where $\mathbf{S}$ is the Katz similarity matrix defined as
@@
Our experimental results on the impact of edge removal, generated through the HOPE embedding method, are presented in Figs. 8 and 9. For the HOPE embedding method, our results suggest that using a 32-dimensional embedding (i.e., subfigures (b) and (d) in the right panel) brings higher ECS scores than a 16-dimensional one (i.e., subfigures (a) and (c) in the left panel). We can also see that HOPE has higher ECS scores with smaller graphs. Analysis of heatmaps in Fig. 10 reveals that the HOPE method tends to preserve effectively higher ECS scores for smaller networks and higher embedding dimensions.
Figure 8: Mean ECS using HOPE method over the percentage of selected nodes on LFR benchmark graphs with $1,000$ nodes. (a) and (b) use random node selection while (c) and (d) use targeted node selection.
Figure 9: Mean ECS using HOPE method over the percentage of selected nodes on LFR benchmark graphs with $10,000$ nodes. (a) and (b) use random node selection while (c) and (d) use targeted node selection.
@@ III.1.4 Modularized non-negative matrix factorization (M-NMF)
The following parameter values are used in our experiments: $\lambda=0.2$, $\alpha=0.05$, $\beta=0.05$, $\eta=5.0$, and number of iterations $N=200$. Our experimental results on the impact of edge removal, generated through the M-NMF embedding method, are reported in Figs. 11 and 12. From these figures, we can see that M-NMF produces higher ECS scores with smaller graphs. Reviewing heatmaps in Fig. 13, M-NMF produces higher ECS scores for strongly clustered networks, but a clear pattern is not discernible. Notably, from Fig. 13(b), M-NMF produces lower ECS scores in cases where $\mu=0.5$. Recall that M-NMF tries to preserve the community structure of networks and therefore is strongly affected by perturbations when the initial community structure is weak.
Figure 11: Mean ECS using M-NMF method over the percentage of selected nodes on LFR benchmark graphs with $1,000$ nodes. (a) and (b) use random node selection while (c) and (d) use targeted node selection.
Figure 12: Mean ECS using M-NMF method over the percentage of selected nodes on LFR benchmark graphs with $10,000$ nodes. (a) and (b) use random node selection while (c) and (d) use targeted node selection.
@@
We derive the following insights for matrix factorization methods. In a nutshell, we observe that ECS scores consistently decrease as we progressively select more nodes and delete their adjacent edges and the networks become sparser. This aligns with our intuition that the community structure is more significantly disrupted when a greater number of edges are removed. From previous figures, it is evident that, in general, for smaller values of $\mu$, the ECS similarity score tends to remain higher for varying degrees of perturbations. This suggests that the graph embedding methods tend to produce higher ECS scores when the original LFR benchmark graph exhibits a stronger partition. Recall that $\mu$ signifies the fraction of nodes sharing edges across different communities. Thus, a lower $\mu$ value implies a higher ratio of internal to external edges, usually resulting in a more pronounced partition in the LFR benchmark graph. Notably, when utilizing targeted node selection, the ECS exhibits a more pronounced and rapid decline compared to random selection. For example, for LFR network with $1,000$ nodes and $\mu=0.5$, when perturbing the network at $50\%$ and using LE with $16$-dimensional embedding, we notice that targeted node selection decreases ECS from $0.67$ to $0.14$ (a decrease of $79.1\%$, see Fig. 2(c)), but ECS decreases from $0.67$ to $0.35$ for random node selection (a decrease of $47.8\%$, see Fig. 2(a)). This observation is comprehensible, as targeted node selection based on betweenness centrality tend to dismantle the network’s community structure more swiftly.
For smaller real-world graphs containing a few thousand vertices, to obtain more robust partitions, the LLE method is suggested with a modest embedding dimension, typically on the order of tens, such as 16 dimensions in our experiments. Alternatively, when the community partition is stronger, the HOPE embedding method with a higher embedding dimension produces more robust partitions. In the case of larger real-world graphs with a weaker community partition, to obtain more robust community partitions, it is suggested to use the LE embedding method with a moderately higher embedding dimension, such as $32$ or $64$ dimensions, as utilized in our experiments. When dealing with larger real-world graphs characterized by a stronger community structure, the LE embedding method with a lower embedding dimension produces more robust community partitions. Alternatively, achieving comparable results can also be accomplished by employing the LLE embedding method with a lower embedding dimension.

### v2

@@ I Introduction
Our findings indicate that incremental selection of nodes and subsequent removal of their adjacent edges consistently reduce ECS. However, this impact tends to be more noticeable across graph embedding methods, but node2vec and LLE, and varies across network sizes, the initial community structure strength, and the perturbation type. This aligns with the expectation that removing a greater number of edges results in a more substantial disruption of the community structure. Notably, targeted node selection leads to a more pronounced and accelerated decline in ECS compared to random node selection. This is rationalized by the fact that adversarial attacks tend to swiftly dismantle the network’s community structure. Among the seven network embedding methods considered, node2vec and LLE consistently prove to be more robust for community detection within their respective graph embedding families; these two methods provide the highest ECS similarity scores. Overall, node2vec is the most robust graph embedding method for community detection. Our results generally concur with those of Kojaku et al. [16], which found that node2vec learns more consistent community structures with networks of heterogeneous degree distributions and size.
@@ II.2 Edge removal procedure
Second, we discuss targeted edge removal. Targeted edge removal is based on the betweenness centrality of nodes. Betweenness centrality measures the importance of a node in a network based on how often it lies on the shortest paths between other nodes [46]. Nodes with high betweenness centrality usually indicate nodes of vital importance to the network, so removing edges adjacent to these nodes might cripple the functionality of the network [47]. This approach models an adversarial attack in real-world situations [48, 49]. To implement edge removal based on targeted node selection, we first rank the nodes in the network by their betweenness centrality. The node with the lowest betweenness centrality has rank $1$, and the node with the highest betweenness centrality has rank $N$, where $N$ is the number of nodes in the network. We then independently generate $50$ sets of nodes, each containing $p\%$ of nodes in the network. The nodes in each set are selected by randomly sampling nodes with probability proportional to their betweenness centrality rank. We ensure that the removal of edges adjacent to each set of nodes does not disconnect the remaining part of the network, which is required by graph embedding methods. In the event that the remaining part of the network becomes disconnected after the removal of a set of edges, we omit the set of nodes and proceed with re-sampling. After each removal, we measure the similarity between the community structure of remaining network in the embedding space computed through data clustering (i.e., $k$-means; see more details in Sec. II.3) and the ground-truth community structure of the original LFR network using ECS. The overall similarity score for a given $p\%$ is calculated by averaging community similarity scores across all $50$ sets.
In our experiments, the proportion of selected nodes starts from $5\%$ with an increment of $5\%$. The upper bound of proportion depends on the network structure and the node selection method. For example, when we select a network with $N=1,000$, $\mu=0.01$, and use edge removal based on targeted node selection, it is difficult to keep the network connected when deleting edges adjacent to $35\%$ of nodes. In these cases, we use a proportion of selected nodes that is below the threshold, e.g., $5\%$, $10\%$, $15\%$, $20\%$, $25\%$, and $30\%$. In addition, we also perform data clustering on the embeddings of the original LFR networks with no edges removed and compare the results with the ground-truth community structure. This result is recorded as “0% of nodes selected.” We also capture the variation of the measurements by computing standard deviations. As the standard deviations are negligible with respect to the mean, we do not show those results here.
@@ II.4 Community similarity
To measure the similarity between community partitions, we employ ECS [35], computed with the CluSim package [61]. Compared to other clustering similarity metrics such as the normalized mutual information (NMI) [62], ECS is a most robust metric for comparing clusters and better addresses challenges such as biases in randomized membership, skewed cluster sizes, and the problem of matching [35]. We use default parameter values in ECS computation. ECS yields values in the range from 00 to $1$, where a higher value indicates a greater similarity between partitions.
@@ III Results and discussion
In the subsequent line figures, unless otherwise stated (in Figs. 16, 24, 25 and 26), subfigures in the top row, i.e., (a) and (b), correspond to random node selection. More specifically, subfigure (a) corresponds to experimental results on small networks with $1,000$ nodes and subfigure (b) depicts experimental results on large networks with $10,000$ nodes. Similarly, subfigures in the bottom row, i.e., (c) and (d), correspond to targeted node selection, for small and large networks respectively.In the body of this paper, we present line figures using $16$-dimensional embedding while line figures with $32$-dimensional embedding are shown in Appendix A. We focus only in these two dimensions as smaller embedding dimensions seem to be enough to provide an accurate low-dimensional vector representation of networks [15, 16]. Recall that a higher ECS value indicates a greater similarity between network community partitions obtained from the LFR benchmark graphs and the embeddings. Thus, when analyzing line figures, higher ECS values and therefore higher positions of line curves indicate a more robust graph embedding method for community detection.
@@
Finally, Sec. III.4 presents experimental results to evaluate the robustness of our community detection methods on two real-world networks with labeled communities: the email-EU-core network [63] and the AS network [7]. These networks exhibit similar characteristics to synthetic counterparts, specifically power-law degree and community size distributions as observed in LFR benchmark graphs. Our experimental methodology and insights for real-world networks mirror findings for synthetic counterparts.
@@ III.1 Matrix factorization methods
Matrix-based methods project network nodes into a Euclidean space via eigenvectors following a similar idea as in spectral clustering [64]. In this family of methods, we include Laplacian eigenmap (LE) [1, 27], locally linear embedding (LLE) [14, 28], higher-order preserving embedding (HOPE) [1, 29], and modularized non-negative matrix factorization (M-NMF) [14, 19]. We now detail results in each of them.
@@ III.1.1 Laplacian eigenmap (LE)
In Fig. 2, the impact of edge removal based on different node selection strategies is demonstrated. The effects of the initial community partition’s strength (with stronger partitions experiencing less perturbation impact, i.e., lower $\mu$ values) and the type of perturbation (targeted node selection having a significant influence on clusters) are observed. Specifically, for networks with strong community partition (i.e., $\mu=0.01$ or $\mu=0.1$), we notice that embedding dimension tends to have a stronger effect on the decay rate of similarity, despite node selection strategy. For instance, in Fig. 2(a), with an embedding dimension of $16$ under random node selection, the curve corresponding to $\mu=0.1$ remains consistently stable around an ECS value of approximately $1.0$ until $60\%$ of nodes are selected. Subsequently, between $60\%$ and $85\%$ of nodes selected, the curve gradually and steadily decreases from $1.0$ to $0.84$ (i.e., $16\%$ decrease). In contrast, depicted in Fig. 17(a) (see Appendix A) using $32$-dimensional embedding, the curve representing $\mu=0.1$ remains stable around $1.0$ only until $25\%$ of node are selected. Then, between $25\%$ and $85\%$ of nodes selected, the curve decreases more sharply from $1.0$ to $0.53$ (i.e., $47\%$ decrease).
Fig. 3 helps us to better understand the effect of embedding dimension and network size. We notice that for strongly clustered networks (i.e., $\mu=0.01$), a lower embedding dimension produces higher ECS. In contrast, for weakly clustered networks (i.e., $\mu=0.5$), lower embedding dimensions produce higher ECS for networks with small sizes and higher embedding dimensions result in higher ECS for networks with larger sizes. In both scenarios represented by Fig. 3(a) and Fig. 3(b), it is observed that, for a given fixed embedding dimension, the data clustering method employing LE yields higher ECS scores for larger networks.
Figure 2: Mean ECS using LE method over the percentage of selected nodes on LFR benchmark graphs with $16$-dimensional embeddings. Random node selection is used in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) are for networks with $1,000$ nodes, and (b) and (d) are for networks with $10,000$ nodes.
@@ III.1.2 Locally linear embedding (LLE)
Figure 4: Mean ECS using LLE method over the percentage of selected nodes on LFR benchmark graphs with $16$-dimensional embeddings. (a) and (b) apply random node selection, whereas (c) and (d) use targeted selection. Networks in (a) and (c) have $1,000$ nodes, and those in (b) and (d) have $10,000$ nodes.
The insights drawn from Figs. 4, 18 and 5 for LLE align closely with the insights derived for LE. In fact, the curves in Fig. 4 are slightly higher in position than those in Fig. 2, indicating that LLE outperforms LE. For example, the data points on the green curve (corresponding to $\mu=0.4$) from LLE in Fig. 4(a) have higher ECS scores over LE in Fig. 2(a) for a given percentage of selected nodes.
@@ III.1.3 Higher-order preserving embedding (HOPE)
Higher-order preserving embedding (HOPE) is designed to maintain the similarity between nodes [1, 29]. The objective function being minimized is $\bigl{|}\mathbf{S}-\mathbf{x}\mathbf{x}^{T}\bigr{|}$, where $\mathbf{S}$ is the Katz similarity matrix defined as
@@
Our experimental results on the impact of edge removal, generated through the HOPE method with embedding dimension $16$, are presented in Fig. 6. For the HOPE embedding method, our results suggest that using a 32-dimensional embedding (i.e., Fig. 19 in Appendix A) brings higher ECS scores than a 16-dimensional one (i.e., Fig. 6). For example, the data points on the green curve (corresponding to $\mu=0.4$) in Fig. 19(a) have higher ECS scores than those data points on the green curve (corresponding to $\mu=0.4$) in Fig. 6(a) at respective percentage of selected nodes. We can also see that HOPE has higher ECS scores with smaller graphs. Analysis of heatmaps in Fig. 7 reveals that the HOPE method tends to preserve effectively higher ECS scores for smaller networks and higher embedding dimensions.
Figure 6: Mean ECS using HOPE method over the percentage of selected nodes on LFR benchmark graphs with $16$-dimensional embeddings. While (a) and (b) use random node selection, (c) and (d) rely on targeted selection. Networks in (a) and (c) contain $1,000$ nodes, while those in (b) and (d) consist of $10,000$ nodes.
@@ III.1.4 Modularized non-negative matrix factorization (M-NMF)
The following parameter values are used in our experiments: $\lambda=0.2$, $\alpha=0.05$, $\beta=0.05$, $\eta=5.0$, and number of iterations $N=200$.
Our experimental results on the impact of edge removal, generated through the M-NMF embedding method with dimension $16$, are reported in Fig. 8. From these figures, we can see that M-NMF produces higher ECS scores with smaller graphs. Our experimental results using $32$-embedding are collected in Fig. 20 (see Appendix A). In Appendix B, we compared these two embedding dimensions for M-NMF. Reviewing heatmaps in Fig. 9, M-NMF produces higher ECS scores for strongly clustered networks, but a clear pattern is not discernible. Notably, from Fig. 9(b), M-NMF produces lower ECS scores in cases where $\mu=0.5$. Recall that M-NMF tries to preserve the community structure of networks and therefore is strongly affected by perturbations when the initial community structure is weak.
Figure 8: Mean ECS using M-NMF method over the percentage of selected nodes on LFR benchmark graphs with $16$-dimensional embeddings. (a) and (b) implement random node selection, while (c) and (d) apply targeted selection; (a) and (c) correspond to networks with $1,000$ nodes, and (b) and (d) to networks with $10,000$ nodes.
@@
III.1.5 Summary of matrix factorization methods
We derive the following insights for matrix factorization methods. In a nutshell, we observe that ECS scores consistently decrease as we progressively select more nodes and delete their adjacent edges and the networks become sparser. This aligns with our intuition that the community structure is more significantly disrupted when a greater number of edges are removed. From previous figures, it is evident that, in general, for smaller values of $\mu$, the ECS similarity score tends to remain higher for varying degrees of perturbations. This suggests that the graph embedding methods tend to produce higher ECS scores when the original LFR benchmark graph exhibits a stronger partition. Recall that $\mu$ signifies the fraction of nodes sharing edges across different communities. Thus, a lower $\mu$ value implies a higher ratio of internal to external edges, usually resulting in a more pronounced partition in the LFR benchmark graph. Notably, when utilizing targeted node selection, the ECS exhibits a more pronounced and rapid decline compared to random selection. For example, for LFR network with $1,000$ nodes and $\mu=0.5$, when perturbing the network at $50\%$ and using LE with $16$-dimensional embedding, we notice that targeted node selection decreases ECS from $0.67$ to $0.14$ (a decrease of $79.1\%$, see Fig. 2(c)), but ECS decreases from $0.67$ to $0.35$ for random node selection (a decrease of $47.8\%$, see Fig. 2(a)). This observation is comprehensible, as targeted node selection based on betweenness centrality tends to dismantle the network’s community structure more swiftly.
For smaller real-world graphs containing a few thousand vertices, to obtain more robust partitions, the LLE method is suggested with a modest embedding dimension, typically on the order of tens, such as 16 dimensions; Alternatively, when the community partition is stronger, the HOPE embedding method with a higher embedding dimension produces comparably robust partitions. In the case of larger real-world graphs with a weaker community partition, to obtain more robust community partitions, it is suggested to use the LE embedding method with a moderately higher embedding dimension, such as $32$ or $64$ dimensions as utilized in our experiments (see Fig. 17 in Appendix A). Alternatively, LLE with a higher embedding dimension produces partitions that are similarly robust but marginally inferior. When dealing with larger real-world graphs characterized by a stronger community structure, the LLE or LE embedding method with a lower embedding dimension produces more robust community partitions. In summary, LLE is the most robust method among matrix factorization methods. Table 3 summarizes our findings for matrix factorization methods.